OpenStudy (anonymous):

Tommy can paint a fence in 9 hours. If his friend Huck helps they can paint it in 1 hour. How fast could Huck paint the fence by himself?

OpenStudy (anonymous):

1/9 + 1/x = 1 1/x = 1-1/9 = 8/9 x = 9/8 hours.